OpenStudy (anonymous):

What is the value of g(x) - f(x) if f(x) = 2x + 4, g(x) = 3x - 2 and x = 12?

OpenStudy (anonymous):

g(x) becomes 3(12)-2 and f(x) becomes 2(12)+4

OpenStudy (anonymous):

so g(x) = 36-2 and f(x) = 24+4 can you solve from there?